# SIGN: Safety-Aware Image-Goal Navigation for Autonomous Drones via Reinforcement Learning

### v1

@@ V-B1 Training Results
The comparison between SIGN and other SoTA algorithms is presented in Table I. Due to the lack of existing continuous baselines, we extend FGPrompt (a discrete SoTA method) to a continuous policy by modifying its action layer. To ensure a fair comparison, both continuous FGPrompt and SIGN are trained for the same number of steps, as illustrated in Fig. 7. On the Gibson benchmark, SIGN outperforms the continuous baseline by +6.4% in SR and +18.3% in SPL. It is worth noting that FGPrompt experiences performance degradation in the continuous setting, highlighting the challenges of continuous control in ImageNav. In contrast, with the augmentation of the proposed auxiliary tasks, SIGN learns an ImageNav policy more efficiently. Regarding SPL, continuous policies do not surpass discrete ones due to the curved trajectories caused by nonzero angular velocity. In the simulator, discrete actions can generate long-range straight paths, yielding higher path efficiency, especially during turns. However, they rely on global position tracking, which is impractical for indoor navigation.
@@ V-C Reliability Analysis of Safety Module
Obstacle avoidance is guided by the collision predictor $Q_{c}$ that is trained together with the navigation policy. Training data comes from raw depth images in the Habitat simulator. To further evaluate its robustness and joint performance with the PX4 flight control system, repeated obstacle avoidance experiments were conducted in Gazebo simulation with pole obstacles. The simulation environment is provided by XTDrone [35], as shown in Fig. 9(a). Gazebo simulator can be seamlessly integrated with PX4 to closely approximate real-life flight dynamics.
@@
In our experiments, depth perception is limited to $3~\mathrm{m}$, with the normalized depth vector $s^{d}_{i}\in[0,1]$ corresponding to a real-world range of $0,3]\mathrm{m}$. Accordingly, the hyperparameter $\beta=0.3$ in Eq.(7) establishes a $0.9~\mathrm{m}$ safety margin, chosen based on the maximal flight speed. Each depth image is cropped to the central $20\%$ horizontal strip, divided into 16 blocks, and the minimum value from each block forms $s^{d}\in\mathbb{R}^{16}$, which is then fed to $Q_{c}$. In Eq. (10), the collision probability threshold is set to $d_{c}=0.5$, corresponding to the Sigmoid function’s decision boundary, and actions are iteratively corrected during navigation to ensure $Q_{c}<d_{c}$.
@@
$\psi(a_{k})=a_{k}-\Delta a=(v_{lin}-\Delta_{lin},v_{ang}-D\cdot\Delta_{ang}),$ (13)
where $D\in\{-1,1\}$ denotes a precomputed safer direction based on the current depth vector $s^{d}$. For a normalized $v_{lin},v_{ang}\in[-1,1]$, we set $\Delta_{lin}=0.3,\Delta_{ang}=0.3$ for both simulations and real-life experiments. To ensure real-time inference, we limit the number of corrections to a maximum of five. If the predicted collision probability is still larger than $d_{c}$ after five corrections, the forward speed will be reset to 0. Then the drone will rotate in place at a speed of $v_{ang}$ to escape from local dead ends.
@@ V-D1 Hardware Configuration
We conduct experiments on a drone equipped with an onboard computer and a RealSense D435i camera. The RGB images have a resolution of 848 × 480 with a $69^{\circ}\times 42^{\circ}$ FoV, while the depth images share the same resolution but feature an $87^{\circ}\times 58^{\circ}$ FoV. Visual-Inertial Odometry (VIO) is executed onboard; however, only the estimated linear velocity and attitude are used for reference tracking control in the PX4 autopilot. The live image stream is transmitted via 2.4 GHz Wi-Fi to an RTX 3060 laptop, with a latency of approximately 10 ms. The velocity control commands $u=[v_{lin},v_{ang}]$ are converted from body frame to global frame commands $u^{\prime}=[v_{x},v_{y},v_{ang}]$ using the yaw angle estimated by VIO, and then sent to the flight controller. The planning loop runs at approximately 20 Hz.
In Fig. 10, real-time velocity commands and the predicted collision probability $Q_{c}$ are displayed in the upper-left corner of the surrounding RGB and depth images, where a positive angular velocity indicates a left turn. The navigation model explores the environment and approaches areas visually similar to the goal image. Fig. 11 shows the flight trajectory along with the predicted collision probabilities at each waypoint.
@@ VI Discussions and Conclusions
In this work, we present SIGN, a novel end-to-end reinforcement learning framework for ImageNav tasks on drones, demonstrating successful sim-to-real transfer. By integrating a visual navigation policy with a safety module, SIGN enables image-goal navigation with reliable collision avoidance. The navigation performance is thoroughly evaluated on three widely used Habitat benchmarks, while the safety module is extensively tested in the Gazebo simulator, which closely approximates real-world flight conditions. Overall, SIGN provides a practical and effective solution for deploying ImageNav on drones.
Our real-world experiments reveal potential failure cases in maze-like environments with visually similar features. When the target image has never been observed, the navigation policy may become stuck in loops, repeatedly exploring the same region without progress. Additionally, due to payload and onboard computational limitations, model inference was not performed onboard in these experiments. In the future, we plan to deploy the algorithms on an NVIDIA Orin platform, enabling planning frequencies of up to 180 Hz during testing.

### v2

@@ V-B1 Training Results
The comparison between SIGN and other SoTA algorithms is presented in Table I. Due to the lack of existing continuous baselines, we extend FGPrompt (a discrete SoTA method) to a continuous policy by modifying its action layer. To ensure a fair comparison, both continuous FGPrompt and SIGN are trained for the same number of steps, as illustrated in Fig. 6. On the Gibson benchmark, SIGN outperforms the continuous baseline by +6.4% in SR and +18.3% in SPL. With the augmentation of the proposed auxiliary tasks, SIGN demonstrates better training efficiency than the baseline.
The performance gap between discrete and continuous policies arises from the increased task complexity. Specifically, discrete policies simplify system dynamics by updating the agent’s pose through teleportation. In contrast, continuous velocity control follows first-order integral dynamics $(\dot{p}=v)$, requiring the agent to reason over continuous transitions and jointly optimize linear and angular velocities. As a result, training continuous-control policies is more challenging for RL, especially for early exploration and policy convergence. The significant performance degradation of FGPrompt when transferred from discrete to continuous actions further highlights the challenges of continuous control in ImageNav.
@@ V-C Reliability Analysis of Safety Module
Obstacle avoidance is guided by the collision predictor $Q_{c}$ that is trained together with the navigation policy. To further evaluate its robustness and joint performance with the PX4 flight control system, repeated obstacle avoidance experiments were conducted in Gazebo with pole obstacles. The simulation environment is provided by XTDrone [xtdrone], as shown in Fig. 8(a). Gazebo simulator can be integrated with PX4 to closely approximate real-life flight dynamics.
@@
In our experiments, depth perception is limited to $3~\mathrm{m}$, and we denote the normalized depth by $s_{t}^{d}$ with $0\leq s_{t}^{d}(i)\leq 1$, corresponding to a real-world range of $0,3]~\mathrm{m}$. Accordingly, the hyperparameter $\beta=0.3$ in Eq.(7) establishes a $0.9~\mathrm{m}$ safety margin, chosen based on the maximal flight speed. Each depth image is cropped to the central $20\%$ horizontal strip, divided into 16 blocks, and the minimum value from each block forms $s_{t}^{d}\in\mathbb{R}^{16}$, which is then fed to $Q_{c}$. In Eq. (10), the collision probability threshold is set to $d_{c}=0.5$, corresponding to the Sigmoid function’s decision boundary, and actions are iteratively corrected to ensure $Q_{c}<d_{c}$.
@@
$\varphi(a_{k})=a_{k}-\Delta a=(v_{lin}-\Delta_{lin},v_{ang}-D\cdot\Delta_{ang}),$ (13)
where $D\in\{-1,1\}$ denotes a precomputed safer direction based on the current depth vector $s_{t}^{d}$. For a normalized $v_{lin},v_{ang}\in[-1,1]$, we set $\Delta_{lin}=0.3,\Delta_{ang}=0.3$ for both simulations and real-life experiments. To ensure real-time inference, we limit the number of corrections to a maximum of five. If the predicted collision probability is still larger than $d_{c}$ after five corrections, the forward speed will be reset to 0. Then the drone will rotate in place at a speed of $v_{ang}$ to escape from local dead ends.
@@ V-D1 Hardware Configuration
We conduct experiments on a drone equipped with an onboard computer and a RealSense D435i camera. The RGB images have a resolution of $848\times 480$ with a $69^{\circ}\times 42^{\circ}$ FoV, while the depth images share the same resolution but feature an $87^{\circ}\times 58^{\circ}$ FoV. Visual-Inertial Odometry (VIO) is executed onboard, but only the estimated linear velocity and attitude are used for reference tracking control in the PX4 autopilot. The live image stream is transmitted via 2.4 GHz Wi-Fi to an RTX 3060 laptop, with a latency of approximately 10 ms. The velocity control commands $u=[v_{lin},v_{ang}]$ are converted from body frame to global frame commands $u^{\prime}=[v_{x},v_{y},v_{ang}]$ using the yaw angle estimated by VIO, and then sent to the flight controller. The planning loop runs at approximately 20 Hz.
In Fig. 9, real-time velocity commands and the predicted collision probability $Q_{c}$ are displayed in the upper-left corner of RGB and depth images. The navigation model explores the environment and approaches areas visually similar to the goal image. Fig. 10 shows the flight trajectory along with the predicted collision probabilities at each waypoint.
@@ VI Discussions and Conclusions
In this work, we present SIGN, a novel end-to-end RL framework for ImageNav tasks on drones, demonstrating successful sim-to-real transfer. By integrating a visual navigation policy with a safety module, SIGN enables image-goal navigation with reliable collision avoidance. The navigation performance is thoroughly evaluated on three Habitat benchmarks, while the safety module is extensively tested in the Gazebo simulator. Overall, SIGN provides a practical and effective solution for deploying ImageNav on drones.
Our real-world experiments reveal potential failure cases in maze-like environments with visually similar features. When the target image has never been observed, the navigation policy may become stuck in loops, repeatedly exploring the same region without progress. In future work, we plan to deploy SIGN onboard the Jetson Orin NX platform with a RealSense camera. In preliminary tests, SIGN took 1.1 GB of GPU memory and the inference frequency reached 30 Hz when VIO was running simultaneously. The CPU utilization was about 35% and the power consumption was 12W.